calculate the time required to deposit 56g of silver from a silver nitrate(AgNO3)solution using a current of 4.5amper


1
Expert's answer
2022-05-26T15:19:03-0400

mass of Ag deposited = 56g 

RAM of Ag = 107.9 g mol-1 (from Periodic Table) 

Moles= (Ag) 56 ÷ 107.9 = 0.519 mole = n(e)

n(e) = 0.519 mol 

F = 96,500 C mol-1

Q = 0.519 x 96,500 = 50,083.5 C

Q = 50,083.5 C  I = 4.5 A 

 

t = Q/I

t = 50,083.5 ÷ 4.5 = 11,129.67 seconds 

t = 11,129.67 ÷ 3600 = 3.1 hours